Calculate the temporary and permanent hardness of water sample containing Mg(HCO3)2= 

46.3mg/L, Ca(HCO3)2= 246.2mg/L, MgCl2= 88.5mg/L, CaSO4=243.6mg/L) in terms of 

CaCO3 equivalents.


1
Expert's answer
2022-01-10T17:02:04-0500

mole of Ca(HCO3)2=246.2×10-3/162

= 1.52×10-3

Mole of CaSO4=243.6×103/136

= 1.79×103

Total of Ca = 3.3×103

"mass of CaCO \n3\n\u200b\t\n =3.3\u00d710^ \n{\u22123}\n \u00d7100=0.33g"

"ppm (permanent hardness) =\frac{ 6.33}{
1000}
 ×10^ 
6
 =200ppm"